1. 3 ducks and two ducklings weigh 32 kg. 4 ducks and 3 ducklings weigh 44 kg. All ducks weigh the same all duckling weigh the same. what is the weight of 2 duck and 1 duckling??
2. Sean has 1 less than twice the number of stickers that jason has. Rick has 53 more stickers than jason has. If sean gives jason 25 stickers. Sean and jason will have the same number of stickers. What is the total number of stickers the 3 boys have?

3. Conrad has been weighing some metal bars. He discovered that 3 silver bars are as heavy as 4 aluminum bars and 5 gold bars weigh as much as 9 silver bars. If he had to weigh 11 gold bars and 24 aluminum bars, Which will be lighter?
4.Amelia forgot to check how much money she began the day with.During the day, she spend $4 on a file, Withdrew $100 from the ATM , got her lunch for $6 and bought 4 Mp3 Pouches at 2 for $12. at the end of the day she had $75. How much did she start the day with?
5. Susan has 4 times as much money as daniel. After giving some money to daniel susan has thrice as much as daniel. then susan spent $40 and gave another $30 to daniel how much does each of them have first?
6. Dan, ethan , felix and gideon have a toal of P197.00 in their wallets. Dan has twice the amount of money felix has. Ethan and gideon together have P152.00 and the total maount of money felix and ethan has is P97.00. How much money does each boy have?
7.Harry paid $28.20 for 2 pieces of rubik's cube and 4 bottles of drinks. Perry paid $52.10 for 5 peices of rubik's cube and 2 bottles of drink. How much does each cost?
8. A number with one decimal place is added to 7. the sum is multiplied by 7 and the product is subtracted by 7 . the difference is divided by 7 , and the result is 9.5 what is the decimal number?
9.Mrs. fields has 7 sons. each son is 3.5 years older than the next younger brother. the oldest son is 5 times and 1 year as old as the youngest brother.how old is each of her sons?


- Please answer this , i need this exactly tomorrow please try your best...thanks

Answer by Bassam(2)

say duck=x and duckling=y
3 ducks= 3x and 2 ducklings= 2y
equation1 3x+2y= 32
equation2 4x+3y= 44
multiply eq1 by 3, 9x+6y=96
multiply eq2 by 2, 8x+6y=88
now (9x+6y)- (8x-6y)= 96-88
9x+6y-8x-6y= 8
tidy up x=8
substitute x=8 in eq1 3X8+2y=32
24+2y=32
2y=32-24
2y=8
y=4
so weight of 2 ducks and 1 ducklings = 2X8+1X4= 20 Kg.